OpenStudy (anonymous):

Which expression is greater(-4)^2/3 or (-4)^3

OpenStudy (texaschic101):

-4^2/3 = -2.519 -4^3 = -64 -4^2/3 is greater

OpenStudy (anonymous):

thank you soo much youve been a big help

OpenStudy (texaschic101):

no problem....you can thank my calculator..lol